OpenStudy (anonymous):

(-17) + 20 – (9 × 8) ÷ 3 = x

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

(-17)+20-(72/3) =(-17)+20-24 =3-24 =-21

OpenStudy (anonymous):

I hope I helped!

OpenStudy (goldphenoix):

(-17) + 20 – (9 × 8) ÷ 3 = x - You do PEMDAS. (-17) + 20 – (72) ÷ 3 = x (-17) + 20 – 24 = x 3 - 24= x -21= x

OpenStudy (anonymous):

Thank you both

OpenStudy (goldphenoix):

Your welcome :)